# Schema Augmentation for Zero-Shot Domain Adaptation in Dialogue State Tracking

### v1

@@ Abstract
Zero-shot domain adaptation for dialogue state tracking (DST) remains a challenging problem in task-oriented dialogue (TOD) systems, where models must generalize to target domains unseen at training time. Current large language model approaches for zero-shot domain adaptation rely on prompting to introduce knowledge pertaining to the target domains. However, their efficacy strongly depends on prompt engineering, as well as the zero-shot ability of the underlying language model. In this work, we devise a novel data augmentation approach, Schema Augmentation, that improves the zero-shot domain adaptation of language models through fine-tuning. Schema Augmentation is a simple but effective technique that enhances generalization by introducing variations of slot names within the schema provided in the prompt. Experiments on MultiWOZ and SpokenWOZ showed that the proposed approach resulted in a substantial improvement over the baseline, in some experiments achieving over a twofold accuracy gain over unseen domains while maintaining equal or superior performance over all domains.
@@
Christopher Richardson1thanks: Equal contribution, Roshan Sharma2111We re-implemented the LDST prompting method using our models., Neeraj Gaur2, Parisa Haghani2, Anirudh Sundar1, Bhuvana Ramabhadran2 1Georgia Institute of Technology, USA 2Google, U.S.A crichardson8@gatech.edu
@@ 1 Introduction
Dialogue State Tracking (DST) plays a pivotal role in task-oriented dialogue systems by maintaining a structured representation of the user’s goals, intents, and preferences as a conversation progresses. As these systems interface with external APIs, such as booking platforms or food ordering services, accurate DST is crucial for ensuring successful goal-oriented interactions. Traditionally, DST models were domain-specific, relying on large amounts of annotated data for every domain. The rise of powerful multi-task instruction-tuned large language models (LLMs) has enabled zero-shot DST, which does not require task-specific training Feng et al. (2023); Yi et al. (2024); Hosseini-Asl et al. (2020). However, zero-shot DST results have been mixed, and prior work has argued that general purpose models remain limited in their ability to replace specialized systems for this task Heck et al. (2023).
Zero-shot domain adaptation offers a promising middle ground. In this problem, training is allowed, but only some domains are available at training time. This allows for methods that retain the benefits of fine-tuning while addressing the challenge of domain-specific data scarcity. Prior work has investigated this method, but has required slots to be filled one at a time, either by generating slot-specific prompts Aksu et al. (2023) or reformulating slot-filling as QA Li et al. (2021); Lin et al. (2021). However, zero-shot domain adaptation methods for end-to-end DST - modeling the entire dialogue state at once, given the dialogue context - is an area which remains underexplored. To this end, we focus on training methods for improving zero-shot domain adaptation with large language models for end-to-end DST. We develop a novel method using data augmentation techniques, and demonstrate the effectiveness of our method on two popular DST datasets, MultiWOZ 2.2 Zang et al. (2020) and SpokenWOZ Si et al. (2024).
Our contributions are threefold:
@@
    We introduce Schema Augmentation, a data augmentation technique for zero-shot domain adaptation that obtains up to a twofold improvement over a strong baseline.
@@
    We propose a new metric, Target Goal Accuracy, to evaluate domain adaptation performance in task-oriented dialogue systems.
  3. 3.
    We conduct ablation studies to analyze the factors contributing to the effectiveness of our method.
@@ 2 Problem Formulation
The goal of Dialogue State Tracking (DST) is to predict the dialogue state given the dialogue text as input. Dialogue states are constructed from sets of slots $\mathcal{S}$ and values $\mathcal{V}$, with each slot belonging to a particular domain $d$. A dataset $\mathcal{X}=\{(x_{i},y_{i})\}$ consists of dialogues $x_{i}$ and dialogue states $y_{i}$. A dialogue state is a set of $k$ distinct slot/value pairs:
$y_{i}=\{(s_{ik},v_{ik})\}_{k}:s\in\mathcal{S},v\in\mathcal{V},$ (1)
The goal is to learn a function that maps dialogue to dialogue state, $\pi(x)=y$. We also have a set of domains $\mathcal{D}$ and say that a dialogue state belongs to a set of domains, $y\in\mathcal{D}$, if $\exists\;(s_{k},v_{k})\in y:s_{k}\in d$ for some $d\in\mathcal{D}$. For zero-shot domain adaptation, we assume there is a set of target domains $\mathcal{D}_{T}$ from which no data is available during training. Thus we have a target subset of our test data
$\mathcal{X}_{\text{target}}=\{(x,y):y\in\mathcal{D}_{T}\}$ (2)
For each slot $s\in\mathcal{S}$ we are given a description of the slot and a list of possible values it can take. This information, referred to as the schema, defines the structure of slot names, descriptions, and possible values, and is typically provided in the prompt.
3 Evaluation Metrics
For evaluation, we utilize Joint Goal Accuracy (JGA) Henderson et al. (2014), the most commonly used DST metric. The main drawback of JGA is that it looks at all slots and domains, and thus does not measure domain adaptation performance directly. To alleviate this, we introduce a new metric: Target Goal Accuracy (TGA), a sub-goal version of JGA that considers only those slots belonging to the target domains.
  • Joint Goal Accuracy (JGA): The percent of turns in which the entire state is accurately predicted. We include all dialogue states, including empty states, to be comprehensive.
  • Target Goal Accuracy (TGA): The percent of turns in which all slots belonging to the target domains are accurately predicted. We do not include turns that have no target-domain slots active in the ground truth dialogue state, as these empty states would dominate the overall metric.
@@ 3 Evaluation Metrics
For TGA we are only interested in the target subset:
$\mathcal{H}_{T}=\{(x,y)\in\mathcal{X}_{\text{target}}:\;\pi(x)=y\}$ (5)
$\text{TGA}=\frac{|\mathcal{H}_{T}|}{|\mathcal{X}_{\text{target}}|}$ (6)
@@
JGA gives us a standardized measure of overall performance on the DST task, while TGA highlights domain adaptation abilities of each method. Our goal is to improve TGA without degradation of JGA.
4 Our Approach: Schema Augmentation
Figure 1: Examples of domain and slot replacements for both Schema Augmentation types: SSA (blue box) and ESA (green box). A full example of augmented prompts can be found in Appendix
Domain adaptation is often addressed by leveraging target domain data for fine-tuning Dingliwal et al. (2021); Zeng et al. (2021) or as few-shot examples during inference Yang et al. (2023); Chen et al. (2023); Rastogi et al. (2019); Hu et al. (2022). However, reliably annotated data for unseen domains is often unavailable. To address this, we propose to achieve zero-shot adaptation through data augmentation. New domains differ from existing ones in their labeling schema and dialogues. Large language models (LLMs), pre-trained on vast amounts of natural language, are likely robust to variations in dialogue, so we focus on differences in schema.
@@ 4 Our Approach: Schema Augmentation
Our method, Schema Augmentation, generates augmented data by altering domain and slot names to enhance the model’s robustness to schema changes. We introduce two types of Schema Augmentation. First, in Synonym Schema Augmentation (SSA), we replace domain and slot names with synonyms. Second, in Encoding Schema Augmentation (ESA), we replace domains and slots with non-semantic codes, requiring the model to rely on slot descriptions and possible values to predict the dialogue state accurately. Figure 1 shows examples of each method.
@@
For both types of augmentation, we experiment with two variations: single and multi. In single-SSA and single-ESA, we use a static synonym or code replacement for each domain and slot. In multi-SSA and multi-ESA, we randomly select the replacement for each sample from a pre-determined list. Gemini Gemini Team (2023) provides the synonym lists, while we create encodings by appending integers to ‘domain’ or ‘slot.’ In multi-ESA, we ensure non-overlapping encodings for each domain and slot by using disjoint ranges of integers. The replacements used in both cases are provided in Appendix A.
5 Experimental Setup
@@ 5.1 Modeling
We perform dialogue state tracking end-to-end using large language models (LLMs). Building on LDST Feng et al. (2023), our prompts include an instruction, DST schema, and dialogue. We represent dialogue state outputs as textual JSON, and during evaluations we parse the output generations using regular expressions (regex). Any response that does not contain valid JSON is considered incorrect by all metrics. Details of our experimental hyperparameters and compute are included in Appendix A.
We choose state-of-the-art instruction-tuned models from two popular families, Gemma Gemma Team (2024) and Mistral Jiang et al. (2023). We use the variants that are most recent at the time of writing: gemma-2-9b-it and Mistral-7B-Instruct-v0.3.
For comparison with prior work, we follow Feng et al. (2023) and build a prompt consisting of instructions, schema, and dialogue. The schema includes slot names, descriptions, and possible values. See Appendix A for full examples of our prompts.
5.2 Datasets
We conduct our experiments on two open-source dialogue state tracking datasets. MultiWOZ Budzianowski et al. (2018) is a multi-domain task-oriented dialogue dataset comprising annotated dialogues across eight domains including hotel booking, restaurant reservation, and taxi ordering. Each dialogue is annotated with the dialogue state at each turn. To alleviate dataset noise, several updated versions of MultiWOZ have since been released. We use MultiWOZ 2.2 Zang et al. (2020), a popular, high quality version of the original. Additionally, we use SpokenWOZ Si et al. (2024), a spoken dialogue TOD dataset inspired by MultiWOZ. SpokenWOZ dialogues were collected from crowdworkers engaging in spoken conversations and includes text transcriptions from an automatic speech recognition (ASR) system. We perform our experiments on the audio transcriptions.
In order to best simulate the domain adaptation scenario, we select holdout domains that minimize slot overlap with the training domains. To this end, we choose {taxi,train,bus} as our holdout domains. Tables showing slot overlap are included in Appendix A.
6 Results
MultiWOZ 2.2 SpokenWOZ
XID Method Model JGATGAJGATGA%
0 Zero-shot prompting (LDST schema) Mistral 8.1 7.2 10.2 6.8
Gemma-2 13.8 16.5 8.2 0.2
@@ 6 Results
Gemma-2 31.1 0.4 25.4 1.2
[…]
Gemma-2 37.9 18.2 32.4 25.7
[…]
Gemma-2 42.1 31.0 32.5 24.2
[…]
Gemma-2 42.8 34.5 30.6 28.8
[…]
Gemma-2 42.7 36.3 33.7 29.4
[…]
Gemma-2 45.2 40.7 34.1 30.3
Table 1: Results for baselines and schema augmentation (out method). We use Huggingface implementations of gemma-2-9b-it and Mistral-7B-Instruct-v0.3. Each experiment is given an experiment ID (XID) for later reference. All experiments use the LDST schema Feng et al. (2023) except the no schema case (XID 1).
Our results are shown in Table 1. We compare our method to several baselines: (0) Zero-shot prompting using LDST prompt from Feng et al. (2023); (1) Standard fine-tuning with no schema, where the input is only the dialogue text; and (2) Fine-tuning using the LDST schema. These baselines are measured against our four methods on both datasets for both models studied. We observe strong performance across the board with both variations of SSA and ESA for both methods, with the strongest performance being achieved with multi-ESA, which boosts TGA from 18.2% to 40.7% on MultiWOZ with Gemma-2.
7 Ablation Studies
LDSTsSSAmSSAsESAmESA10101010202020203030303040404040TGA (%) vs. ExperimentOriginalShuffled Descriptions/Values
Figure 2: Ablation results for Gemma-2 on MultiWOZ 2.2. Experiment names are abbreviated. LDST=no augmentation; single/multi are shortened to s/m.
Our results in Table 1 raise the following questions:
  1. 1.
    Why does schema augmentation (both SSA and ESA) improve domain adaptation?
  2. 2.
    Why does ESA improve adaptation more than SSA?
To shed light on these questions, we perform ablation studies by shuffling descriptions and possible values at test time and observing the performance impact. This ablation tests the dependence of each method on the information content of the descriptions and values, without introducing confounding variables related to mismatch between train and test samples.
Figure 2 shows the ablation results for Gemma-2 on MultiWOZ 2.2. To answer the Question 1, with description/value shuffling, we observe a greater drop in TGA for Schema Augmentation (min: -14.5%, max: -24.2%) than for the LDST baseline (-7.8%). This suggests that Schema Augmentation encourages the model to pay more attention to the descriptions and values of the slots. This becomes particularly useful for target domain samples for which the model has not seen any training data. Additionally, we observe that TGA drop is greater for the ESA (sESA: -18.5%, mESA: -24.2%) than for SSA (sSSA: -15.7%, mSSA: -14.5%), suggesting that ESA more strongly encourages the model to attend to descriptions and values provided in the schema.

### v2

@@ Abstract
Zero-shot domain adaptation for dialogue state tracking (DST) remains a challenging problem in task-oriented dialogue (TOD) systems, where models must generalize to target domains unseen at training time. Current large language model approaches for zero-shot domain adaptation rely on prompting to introduce knowledge pertaining to the target domains. However, their efficacy strongly depends on prompt engineering, as well as the zero-shot ability of the underlying language model. In this work, we devise a novel data augmentation approach, Schema Augmentation, that improves the zero-shot domain adaptation of language models. Schema Augmentation is a simple but effective technique that enhances generalization by introducing variations of slot names within the schema provided in the prompt. Experiments on MultiWOZ and SpokenWOZ showed that the proposed approach resulted in a substantial improvement over the baseline, in some experiments achieving over a twofold accuracy gain over unseen domains while maintaining equal or superior performance over all domains.
@@
Christopher Richardson1thanks: Equal contribution, Roshan Sharma2111Our reimplementation of Zhao et al. (2022)., Anirudh Sundar1111Our reimplementation of Zhao et al. (2022)., Neeraj Gaur2, Parisa Haghani2, Bhuvana Ramabhadran2 1Georgia Institute of Technology, USA 2Google, U.S.A crichardson8@gatech.edu
@@ 1 Introduction
Figure 1: Examples of domain and slot replacements for both Schema Augmentation types: SSA (blue box) and ESA (green box).
An essential problem for task-oriented dialogue systems is Dialogue State Tracking (DST), the task of extracting a structured respresentation of the dialogue state from of user goals as the conversation progresses (see Table 1). While traditional DST models require extensive domain-specific annotation, instruction-tuned large language models (LLMs) enable zero-shot DST Feng et al. (2023); Yi et al. (2024); Hosseini-Asl et al. (2020). However, their effectiveness remains limited compared to specialized systems Heck et al. (2023).
Zero-shot domain adaptation bridges this gap by allowing training on a subset of domains, improving generalization without domain-specific data Aksu et al. (2023); Li et al. (2021); Lin et al. (2021a). However, little work has explored end-to-end DST—modeling the full dialogue state in a single step. We address this gap by developing a novel data augmentation approach for training LLMs on zero-shot domain adaptation and evaluating it on MultiWOZ 2.1 Zang et al. (2020) and SpokenWOZ Si et al. (2024).
Our main contributions are:
@@
    We introduce Schema Augmentation, a data augmentation method that improves zero-shot domain adaptation, achieving up to a twofold improvement over strong baselines.
@@
    We propose Target Goal Accuracy, a new metric for evaluating domain adaptation in task-oriented dialogue.
2 Related Work
Prior work on zero-shot domain adaptation for dialogue state tracking has focused on the use of semantic slot information to generalize to unseen domains Rastogi et al. (2020). T5DST Lin et al. (2021b) utilized the slot type for more descriptive semantic slots and showed that using slot-type in conjunction with descriptions improved the zero-shot Joint Goal Accuracy. Lee et al. (2021) presented an approach that included natural language descriptions of the schema in the prompt.
Prior work has also explored generalization in DST by reframing the task as question-answering Li et al. (2021); Heck et al. (2024). Other approaches include parameter-efficient methods such as Prompter Aksu et al. (2023), and DualLoRA Luo et al. (2024). Prompter applied prompt-based learning with in-context examples, while DualLoRA separated the adapters used for the prompts and the dialogue context to assist generalizability. D3ST Zhao et al. (2022) relied solely on schema descriptions and represents the current state of the art in zero-shot domain adaptation on MultiWOZ 2.1. In contrast to D3ST that replaces the schema, this work focuses on augmenting the dataset with new samples with modified schemata. Prior work has provided preliminary evidence on the benefits of data augmentation Ma et al. (2019) by back-translating sentences between English and Chinese. This work formalizes the schema augmentation using synonym augmentation from a controlled set of synonyms and an encoding-based schema augmentation that relies on random replacement.
@@ 3 Problem Formulation
Dialogue states are constructed from sets of slots $\mathcal{S}$ and values $\mathcal{V}$, with each slot belonging to a particular set of domains $\mathcal{D}$. A dataset $\mathcal{X}=\{(x,y)\}$ consists of dialogues $x$ and dialogue states $y$. A dialogue state is a set of $K$ distinct slot/value pairs, where $K\geq 0$ and differs for every $(x,y)$:
$y=\{(s_{k},v_{k})\}_{k=1}^{K}:s\in\mathcal{S},v\in\mathcal{V}$ (1)
The goal is to learn a function that maps dialogue to dialogue state, $\pi(x)=y$. When evaluating domain adaptation, we are particularly interested in predicting slots that belong to our target domains, which we will call $\mathcal{D}_{T}$. Thus the dialogue substate of interest $y_{T}\subseteq y$ is:
$y_{T}\overset{\Delta}{=}\{(s,v)\}:s\in\mathcal{D}_{T}$ (2)
Similarly, we denote the target-domain subset of the predicted state as $\pi_{T}(x)$. To most accurately reflect true domain adaptation performance, we look at only those dialogues that have nonempty target-domain substates:
$\mathcal{X}_{T}=\{(x,y_{T})\}:y_{T}\neq\varnothing$ (3)
For each slot $s\in\mathcal{S}$ we are given a description of the slot and a list of possible values it can take. This information, referred to as the schema, defines the structure of slot names, descriptions, and possible values, and is provided in the prompt.
4 Methods
4.1 Target Goal Accuracy
The primary metric used in the DST literature is Joint Goal Accuracy (JGA) Henderson et al. (2014), which looks at all slots and domains, and thus does not measure domain adaptation performance directly. To alleviate this, we introduce a new metric: Target Goal Accuracy (TGA), a sub-goal version of JGA that considers only those slots belonging to the target domains.
Joint Goal Accuracy (JGA):
The fraction of turns in which the entire state is accurately predicted. We include all dialogue states, including empty states, to be comprehensive.
Target Goal Accuracy (TGA):
The fraction of all target-domain turns (turns with at least one slot belonging to the target domains) in which the target-domain substate is accurately predicted. We do not include turns that have no target-domain slots active in the ground truth dialogue state, as these empty states would dominate the overall metric.
@@ 4.1 Target Goal Accuracy
TGA only considers the target subset:
$\mathcal{H}_{T}=\{(x,y_{T})\in\mathcal{X}_{T}:\;\pi_{T}(x)=y_{T}\}$ (6)
$\text{TGA}=\frac{|\mathcal{H}_{T}|}{|\mathcal{X}_{T}|}$ (7)
@@
JGA gives a standardized measure of overall performance on DST, while TGA highlights domain adaptation abilities of each method. Our goal is to improve TGA without degradation of JGA.
Turn Dialogue Dialogue State Model Response JGA TGA
USER: Hello I’d like to book a table for Friday. restaurant-day: friday restaurant-day: friday
1 SYSTEM: I see openings at Chotchkie’s all evening, is that ok? restaurant-time: - restaurant-time: - N/A
taxi-arriveby: - taxi-arriveby: -
USER: Yes, let’s do 19:00. restaurant-day: friday restaurant-day: friday
2 SYSTEM: Great, it’s booked. Can I help you with anything else? restaurant-time: 19:00 restaurant-time: 19:00 N/A
taxi-arriveby: - taxi-arriveby: -
USER: I also need a taxi to pick me up 30 minutes before. restaurant-day: friday restaurant-day: friday
3 SYSTEM: Okay, I scheduled it. restaurant-time: 19:00 restaurant-time: 19:00
taxi-arriveby: 18:30 taxi-arriveby: -
Total 67% 0%
Table 1: Example dialogue illustrating the difference between JGA and TGA for the taxi domain. The last two columns show whether the response triggers a hit or miss for the correpsponding metric, and N/A means that example will not be counted in the metric. TGA better reflects performance on taxi by ignoring samples with empty taxi substates. A model that never attempts to generate taxi slots can potentially still achieve a high JGA.
We are motivated by the limitations of prior work in measuring true domain adaptation, which follows a "leave-one-out" cross-domain adaptation setup Lin et al. (2021a, b). In this setup, one domain is withheld from training, and the model is trained on the remaining domains before being tested on the holdout domain. We argue that this does not capture true domain adaptation due to two key issues: slot overlap and empty substate bias.
Slot overlap occurs when slots are shared across domains, reducing the challenge of encountering unseen slots. Table 5 shows that domains like taxi/train/bus and hotel/restaurant have significant slot overlap, making adaptation easier when similar slots exist in training.
The empty substate bias arises from how prior work evaluates zero-shot domain performance using JGA. The test set is filtered to include only samples labeled with the target domain, and only target domain slots are measured for JGA calculation. However, prior work does not filter out empty slots, leading to cases where the entire substate for the target domain is empty. This is especially problematic for domains appearing later in dialogues, such as taxi, resulting in artificially inflated JGA scores.
Table 1 illustrates this issue with a MultiWOZ-style example. A model with no taxi training never fills taxi slots. When evaluating JGA for taxi, turns 1-2 count as correct because the model "predicted" the empty taxi substate—despite lacking any knowledge of taxi. In real-world scenarios, such "hits" are irrelevant, so we propose TGA to better reflect true domain adaptation. In this example, JGA is 2/3 (67%), while TGA is 0/1 (0%).
4.2 Schema Augmentation
@@
In our method, Schema Augmentation, we generate augmented data by altering domain and slot names to enhance the model’s robustness to schema changes. We explore two types of Schema Augmentation: Synonym Schema Augmentation (SSA), in which we replace domain and slot names with synonyms; and Encoding Schema Augmentation (ESA), in which we replace domains and slots with non-semantic codes, requiring the model to rely on slot descriptions and possible values to predict the dialogue state accurately. Figure 1 illustrates and example of these two types of replacements.
@@
For both types of augmentation, we utilize multiple possible replacements for each slot and domain. We randomly select the replacement for each sample from a pre-determined list. Gemini Gemini Team (2023) provides the synonym lists, while we create encodings by appending integers to ‘domain’ or ‘slot.’ In ESA, we ensure non-overlapping encodings for each domain and slot by using disjoint ranges of integers. Appendix A describes replacements used in both cases.
5 Experiments
@@ 5.1 Modeling
We perform dialogue state tracking end-to-end by fine-tuning gemma-2-9b-it Gemma Team (2024) on prompts constructed from dialogue data paired with dialogue state outputs. We build on LDST Feng et al. (2023) for our prompt construction. Our prompts include an instruction, DST schema, and dialogue, and outputs are represented as textual JSON. Details of our experimental hyperparameters and compute are included in Appendix A, as well as full examples of our prompts.
All of our experiments use Gemma Gemma Team (2024). We utilize the gemma-2-9b-it variant as it is feasible to train on a single GPU, and we use the instruction tuned version in order to enable understanding of language feedback.
5.2 Datasets and Baselines
We conduct our experiments on two open-source dialogue state tracking datasets — MultiWOZ2.1 Eric et al. (2019) and SpokenWOZ Si et al. (2024), a spoken TOD dataset inspired by MultiWOZ. For ablation studies, we perform experiments on MultiWOZ 2.2 Zang et al. (2020), a cleaner update to 2.1 with less noise. A detailed description is provided in A.3.
@@ 5.2 Datasets and Baselines
We compare our approach against several baselines from prior work on zero-shot dialogue state tracking and cross-domain transfer — TransferQA, T5DST, Prompter, DualLoRA, and D3ST. The methods are described in Section 2.
Method $\textbf{JGA}_{\text{taxi}}$ $\textbf{JGA}_{\text{train}}$ TGA
Prior work
TransferQA 61.9 36.7 15.9
T5DST 64.6 35.4 15.7
Prompter 66.3 39.0 20.2
DualLoRA 67.2 42.4 24.2
D3ST 78.4 38.7 25.2
Our approach
SSA 66.7 48.8 31.8
ESA 69.5 50.6 34.9
Table 2: Comparison of our method to prior work on MultiWOZ 2.1. Goal accuracies (%) showing JGA reported in original prior works with the corresponding estimated TGA.
5.3 Main Results
Table 2 shows results on the MultiWOZ 2.1 dataset alongside results from prior work. For direct comparison to our method, we only report results for JGA on taxi and train domains, as well as TGA. Details of computation of TGA from JGA for prior work can be found in the Appendix A.2. Our method shows an improvement of 19.6% TGA (absolute) over the best performing baseline, D3ST Zhao et al. (2022), despite using a model with 20% fewer parameters (9B vs. 11B). There are a few things to note about approaches in prior work that confer an advantage. For one, they do not address the issue of slot overlap. Previous experiment using the leave-one-out approach ignore the slot overlap problem (see Section 4.1 for reference) and train models on many of the same slots that are seen in the target domain. This gives a large advantage, particularly when evaluating on domains like taxi whose slots are all included in domains (train) present in the training data.
Our main results are shown in Table 2. Our best method, ESA, improves TGA over the best performing baseline, D3ST, by 9.7%. It also improves JGA on train by 11.9% and achieves the second highest JGA on taxi, despite having no access to those slots during training time, unlike all the baselines we compare to.
We also report results using Mistral-7B-Instruct-v0.3 Jiang et al. (2023), which is shown in the Appendix.
5.4 SpokenWOZ Results
Method $\textbf{JGA}_{\text{taxi}}$ $\textbf{JGA}_{\text{train}}$ TGA
Prior work
D3ST 11footnotemark: 1 67.1 41.1 21.9
Our approach
SSA 67.6 49.6 28.8
ESA 66.7 50.5 30.3
Table 3: SpokenWOZ results. Comparison of our method to the prior state-of-the-art D3ST. Note that JGA here is overall JGA on the full test set with all domains.
To further assess our method, we also run experiments on SpokenWOZ. To the best of our knowledge, no prior work on cross-domain transfer has been done on SpokenWOZ, so we re-implement the highest performing baseline from MultiWOZ (D3ST, Zhao et al. (2022)) for comparison. The results are shown in Table 3, with our method outperforming D3ST.